# FedHERO: A Federated Learning Approach for Node Classification Task on Heterophilic Graphs

### v1

@@ 1 Introduction
Graph Neural Network (GNN) aims to extract informative patterns from graphs Xia et al. (2021); Wu et al. (2020b); Zhou et al. (2020); Scarselli et al. (2008); Xu et al. (2018a). However, graph data may contain sensitive information about the involved individuals (e.g., a user’s race in a social network Dwork et al. (2012); Beutel et al. (2017)). As a consequence, sharing graph data across different stakeholders is raising concerns due to the potential risk of privacy leakage, which further prevents training GNNs in a centralized manner Cong & Mahdavi (2022); Dai et al. (2022); Mei et al. (2019); Zhu & Zhu (2020). To properly handle the above-mentioned problems, Federated Graph Learning (FGL) has emerged popular as a collaborative learning framework, which enables participants (clients) to derive insights from distributed graph sources while upholding stringent privacy safeguards Liu et al. (2022a); Lou et al. (2021); Zhu et al. (2022); Zhang et al. (2022); Peng et al. (2021); Chen et al. (2021b). A typical scenario of FGL is that each client possesses a single graph (local graph) and seeks to collaboratively train a global GNN model for certain tasks (e.g., node classification tasks Xiao et al. (2022); Bhagat et al. (2011)) without sharing the raw graph data. In such scenarios, current FGL approaches implicitly rely on the assumption of homophilic local graphs Zhu et al. (2020); Ma et al. (2021), i.e., nodes within a client’s graph are more likely to form connections with other nodes of the same class Ma et al. (2019); Tang et al. (2009b). Specifically, when clients hold homophilic graphs, most of the neighbors of each node will have the same class as this center node Ying et al. (2018). The GNNs trained on different clients will thus capture a similar tendency, and aggregating these GNNs usually makes the global model in FGL better learn such a tendency and benefit the generalization ability of the global model Baek et al. (2023); Zhang et al. (2021).
@@
Nevertheless, a critical issue regarding the above-mentioned assumption is that it overlooks the prevalence of heterophilic graphs Pandit et al. (2007); Zheng et al. (2022; 2023); Bo et al. (2021); Yang et al. (2021); Chien et al. (2020); Liu et al. (2023a), where nodes belonging to different classes tend to be connected. For instance, in financial transaction networks, nodes denote bank clients, where node labels are their credit risk and edges represent the financial interactions between them. Such a network can be heterophilic, and we show an example in Figure 1. Specifically, a ‘Moderate Risk’ customer in Bank A may transact with multiple ‘Low Risk’ individuals, whereas in Bank B, the ’Moderate Risk’ customer engages primarily with ’High Risk’ counterparts. When training GNNs to predict customer risk, Bank A’s GNN learns the connection between ’Moderate Risk’ and ’Low Risk,’ while Bank B’s GNN captures the connection between ’Moderate Risk’ and ’High Risk.’ Directly aggregation models could lead to poor performance in predicting ’Moderate Risk’ customers, as they fail to establish a consistent pattern that generalizes well across both banks. Consequently, this mismatch can result in performance degradation for FGL algorithms.
@@ 2 PRELIMINARIES
Federated Learning. The objective of clients in FL is to jointly train models using local data. Let’s consider $M$ data owners, each possessing data exclusively accessible to them, denoted as $\mathcal{D}_{i}=(\mathcal{X}_{i},\mathcal{Y}_{i})$, where $\mathcal{X}_{i}$ represents the data instances and $\mathcal{Y}_{i}$ is the corresponding labels. $N_{i}$ signifies the number of data samples held by client $i$ and $N=\sum_{i}N_{i}$ denotes the total number of data samples available. $\mathcal{L}_{i}$ and $w_{i}$ are the loss function and model parameters of client $i$, respectively. The FL process typically encompasses the following steps McMahan et al. (2017): (i) At $r$-th round, the server selects a subset of clients that participate in the training process. These clients receive the global model from the last round $\overline{w}_{r}$ and initialize the local model parameters as $w_{i}^{r}\leftarrow\overline{w}_{r}$. (ii) Each selected client updates its local model using its own local data to minimize the task loss: $w_{i}^{r+1}\leftarrow w_{i}^{r}-\eta\nabla\mathcal{L}_{i}(w_{i}^{r})$. (iii) The server aggregates the updated local models to obtain an updated global model: $\overline{w}_{r+1}=\sum_{i}\frac{N_{i}}{N}w_{i}^{r+1}$. This process continues iteratively until convergence is achieved.
Preliminary Study. To highlight the unique challenges posed by varying neighbor distribution patterns across clients, we conduct a preliminary study comparing the performance of local training models with the global model produced by FedAvg. Table 1 contrasts results on heterophilic datasets Rozemberczki et al. (2021) (Squirrel and Chameleon) with those on homophilic datasets (PubMed Namata et al. (2012) and Citeseer Sen et al. (2008)). On homophilic graphs, the global model generally shows better performance. In contrast, for heterophilic graphs, the aggregated global model underperforms relative to local training models. This discrepancy arises because clients with heterophilic graphs are more likely to exhibit diverse neighbor distribution patterns, leading locally trained GNNs to capture varied tendencies. As a result, the global model struggles to aggregate and generalize this knowledge, resulting in poor predictions. This preliminary experiment underscores the impact of distinct neighbor distribution patterns on FGL performance, reinforcing the need to address the challenges identified in our study.
@@ 3.1 Graph Structure Learning in FedHERO
Graph structure learning endeavors to comprehend the message-flow pattern throughout the graph Xia et al. (2021). Nevertheless, in FL scenarios where data is diversely distributed and privacy concerns are paramount, directly aggregating all data to train a comprehensive structure learning model becomes unfeasible. To tackle this challenge, we implement a structure learning model for each client, which can learn structural knowledge from local data. By sharing the learned knowledge with other clients, the entire FGL system can acquire a unified, globally optimized structural representation that captures universally applicable patterns across diverse graphs.
@@
The objective of the structure learning model is to generate relevant latent structures that facilitate effective message passing based on the original graph. A common approach considers representing the weight of an edge between two nodes as a distance measure between two end nodes Zhao et al. (2021b); Kazi et al. (2022); Chen et al. (2020). We follow this idea and employ a one-layer GNN to derive node representations $\mathbf{Z}$ from the original graph. The structure learning model can be represented by learning a metric function $\phi(\cdot,\cdot)$ of a pair of representations: $\tilde{a}_{uv}=\phi(\mathbf{z}_{u},\mathbf{z}_{v})$, where $\mathbf{z}_{u}$ is the learned embedding of node $u$ produced by GNN. The metric function $\phi(\cdot,\cdot)$ can assume various forms like cosine distance Nguyen & Bai (2010) and RBF kernel Li et al. (2018b). In this study, we employ the multi-head attention mechanism as the metric function, following the approach outlined by Chen et al. Chen et al. (2020):
@@
This design choice is based on the differentiability of the top-k function, which facilitates the computation of parameter gradients and subsequent model updates in FedHERO. Additionally, we explore an alternative method for latent graph generation. Such a method treats each edge in the latent graph as a Bernoulli random variable. Here, each latent edge $\tilde{\mathbf{A}}_{ij}$ is sampled from a Bernoulli distribution with parameter $\tilde{a}_{ij}$ Wu et al. (2020a); Elinas et al. (2020); Zhao et al. (2023). A comparison of these methods and the impact of varying k values is provided in Appendix C.
@@ 3.2 Optimization of the Dual-channel Model
In existing FGL frameworks Zhang et al. (2021); Baek et al. (2023), the local models primarily employ one or multiple GNNs in sequence. These models initiate the process of generating node representations directly from raw features. However, as Tan et al. Tan et al. (2023) pointed out, directly constructing and sharing these local models across diverse clients can significantly compromise personalized performance due to feature and representation space misalignment. To address this challenge, divide the local model into two channels: $f_{global}$ and $f_{local}$. The global learner $f_{global}$ takes the original input graph and refines the latent graph structure. Through cooperation with other clients, $f_{global}$ detects patterns of universal relevance across heterophilic graphs. On the other hand, the local task learner $f_{local}$ extracts valuable client-specific information from raw node features and structures within the local graph. This information includes biased structures or diverse feature distributions tailored for follow-up tasks.
In a client’s local graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, we initiate the process by utilizing a Multi-Layer Perception (MLP) $f_{0}$ to extract information from the node feature matrix $\mathbf{X}$ into node embeddings $\mathbf{Z}^{0}$. Next, we feed the node embeddings $\mathbf{Z}^{0}$, the latent graph $\tilde{\mathbf{A}}$, and the original graph structure $\mathbf{A}$ into an L-layer GNN. The middle part of Figure 2 depicts how FedHERO leverages information from both the local graph structure $\mathbf{A}$ and the structure $\tilde{\mathbf{A}}$ from the structure learner to obtain node representations. In the $f_{global}$ channel (models in yellow boxes in Figure 2), the l-th layer of the GNN (denoted as $f_{g}^{l}$) utilizes the node embeddings from the previous layer $\mathbf{Z}^{l-1}$ and the latent graph $\tilde{\mathbf{A}}$ to generate hidden node embeddings that reveal patterns that hold universally across different graphs, denoted as $\mathbf{H}^{l}$. While feature propagation on the $\tilde{\mathbf{A}}$ could potentially yield enhanced node embeddings that align with global patterns, it is important to note that the original graph structures also contain valuable information, such as effective inductive bias Bronstein et al. (2017). Therefore, we leverage the information within $\mathbf{A}$. It is combined with the output from the last layer, $\mathbf{Z}^{l-1}$, and fed into the l-th layer of the GNN (denoted as $f_{loc}^{l}$) within the $f_{local}$ channel (models in blue boxes in Figure 2). This process generates hidden node embeddings that adhere to the message-passing patterns within the local graph, resulting in $\mathbf{E}^{l}$. These embeddings play a crucial role in computing the layer-wise updates for node representations:
@@
Recent studies show that merging intermediate node representations from multi-hops enables the GNN to integrate nodes of the same class Zhu et al. (2020); Abu-El-Haija et al. (2019). In addition, encoding the features of the ego node separately from the aggregated neighbor representations enhances node embedding learning Zhu et al. (2020); Suresh et al. (2021); Yan et al. (2022). Thus, we merge the intermediate node representations generated by each layer of the dual-channel GNNs with the ego feature $\mathbf{X}$: $\mathbf{Z}^{out}=\textrm{COMB}(\mathbf{X},\mathbf{Z}^{0},...,\mathbf{Z}^{L}).$ The COMB function can be implemented in several ways, including max-pooling, concatenation, or LSTM-attention Xu et al. (2018b). In our approach, we specifically use concatenation for the COMB function. This choice is based on the observation that concatenation consistently outperforms other methods in similar contexts Xu et al. (2018b); Zhu et al. (2020).
@@
where $||\cdot||_{F}$ is the Frobenius norm. The first term in Eq. (7) evaluates the smoothness of the latent graph Belkin & Niyogi (2001), operating under the assumption that nodes with smoother features are more likely to be connected. The second term in Eq. (7) serves the purpose of preventing excessively large node degrees Kalofolias (2016). The hyperparameters $\lambda$ and $\mu$ govern the magnitude of the regularization effects. Hence, in conjunction with the cross-entropy loss $\mathcal{L}_{ce}$, we fine-tune the local model using the following loss function:
@@ 3.3 Model Aggregation in FedHERO
The right side of Figure 2 illustrates the process of clients sharing local models and receiving updated models from the server. The model aggregation process in FedHERO resembles that of FedAvg McMahan et al. (2017). However, in the FedHERO framework, the server only receives and aggregates models from the $f_{global}$ channel of clients instead of the whole model. This sharing mechanism offers two key advantages for FedHERO. First, the structure learning model in the global channel generates latent graphs from node embeddings with relatively low dependence on the original graph structure. By sharing this model among clients, the global channel GNNs learn similar patterns from the generated latent graphs. Aggregating these aligned models enhances the effectiveness of FGL, leading to improved overall performance. Second, retaining the model $f_{local}$—which focuses on learning the message passing pattern within the local graph—as private and training it locally allows the local model to capture client-specific information, ultimately contributing to subsequent tasks. Let $w_{g}^{i}$ denote the parameters of the $f_{global}$ in client $i$. The server does the aggregation: $\overline{w}_{g}=\sum_{i}\frac{N_{i}}{N}w_{g}^{i}$ where $N_{i}$ represents the number of nodes in the client $i$’s graph, and $N$ denotes the total number of nodes in the graph data of all clients. Subsequently, the server transmits $\overline{w}_{g}$ to the clients. The clients then update their model in the $f_{global}$ channel with $\overline{w}_{g}$ and commence local training for the ensuing round.

### v2

@@ 1 Introduction
Graph Neural Network (GNN) aims to extract informative patterns from graphs (Xia et al., 2021; Wu et al., 2020b; Zhou et al., 2020; Scarselli et al., 2008; Xu et al., 2018a). However, graph data may contain sensitive information about the involved individuals (e.g., a user’s race in a social network (Dwork et al., 2012; Beutel et al., 2017)). As a consequence, sharing graph data across different stakeholders is raising concerns due to the potential risk of privacy leakage, which further prevents training GNNs in a centralized manner (Cong & Mahdavi, 2022; Dai et al., 2022; Mei et al., 2019; Zhu & Zhu, 2020). To properly handle the above-mentioned problems, Federated Graph Learning (FGL) has emerged as a popular collaborative learning framework, which enables participants (clients) to derive insights from distributed graph sources while upholding stringent privacy safeguards (Liu et al., 2022a; Lou et al., 2021; Zhu et al., 2022; Zhang et al., 2022; Peng et al., 2021; Chen et al., 2021b). A typical scenario of FGL is that each client possesses a single graph (local graph) and seeks to collaboratively train a global GNN model for certain tasks (e.g., node classification tasks (Xiao et al., 2022; Bhagat et al., 2011)) without sharing the raw graph data. In such scenarios, current FGL approaches implicitly rely on the assumption of homophilic local graphs (Zhu et al., 2020; Ma et al., 2021), i.e., nodes within a client’s graph are more likely to form connections with other nodes of the same class (Ma et al., 2019; Tang et al., 2009b). Specifically, when clients hold homophilic graphs, most of the neighbors of each node will have the same class as this center node (Ying et al., 2018). The GNNs trained on different clients will thus capture a similar tendency, and aggregating these GNNs usually makes the global model in FGL better learn such a tendency and benefit the generalization ability of the global model (Baek et al., 2023; Zhang et al., 2021).
@@
Nevertheless, a critical issue regarding the above-mentioned assumption is that it overlooks the prevalence of heterophilic graphs (Pandit et al., 2007; Zheng et al., 2022; 2023; Bo et al., 2021; Yang et al., 2021; Chien et al., 2020; Liu et al., 2023a), where nodes belonging to different classes tend to be connected. For instance, in financial transaction networks, nodes denote bank clients, where node labels are their credit risk and edges represent the financial interactions between them. Such a network can be heterophilic, and we show an example in Figure 1. Specifically, a ‘Moderate Risk’ customer in Bank A may transact with multiple ‘Low Risk’ individuals, whereas in Bank B, the ’Moderate Risk’ customer engages primarily with ’High Risk’ counterparts. When training GNNs to predict customer risk, Bank A’s GNN learns the connection between ’Moderate Risk’ and ’Low Risk,’ while Bank B’s GNN captures the connection between ’Moderate Risk’ and ’High Risk.’ Directly aggregation models could lead to poor performance in predicting ’Moderate Risk’ customers, as they fail to establish a consistent pattern that generalizes well across both banks. Consequently, this mismatch can result in performance degradation for FGL algorithms.
@@ 2 PRELIMINARIES
Federated Learning. The objective of clients in FL is to jointly train models using local data. Let’s consider $M$ data owners, each possessing data exclusively accessible to them, denoted as $\mathcal{D}_{i}=(\mathcal{X}_{i},\mathcal{Y}_{i})$, where $\mathcal{X}_{i}$ represents the data instances and $\mathcal{Y}_{i}$ is the corresponding labels. $N_{i}$ signifies the number of data samples held by client $i$ and $N=\sum_{i}N_{i}$ denotes the total number of data samples available. $\mathcal{L}_{i}$ and $w_{i}$ are the loss function and model parameters of client $i$, respectively. The FL process typically encompasses the following steps: (i) At $r$-th round, the server selects a subset of clients that participate in the training process. These clients receive the global model from the last round $\overline{w}_{r}$ and initialize the local model parameters as $w_{i}^{r}\leftarrow\overline{w}_{r}$. (ii) Each selected client updates its local model using its own local data to minimize the task loss: $w_{i}^{r+1}\leftarrow w_{i}^{r}-\eta\nabla\mathcal{L}_{i}(w_{i}^{r})$. (iii) The server aggregates the updated local models to obtain an updated global model: $\overline{w}_{r+1}=\sum_{i}\frac{N_{i}}{N}w_{i}^{r+1}$. This process continues iteratively until convergence is achieved.
Preliminary Study. To highlight the unique challenges posed by varying neighbor distribution patterns across clients, we conduct a preliminary study comparing the performance of local training models with the global model produced by FedAvg. Table 1 contrasts results on heterophilic datasets (Rozemberczki et al., 2021) (Squirrel and Chameleon) with those on homophilic datasets (PubMed (Namata et al., 2012) and Citeseer (Sen et al., 2008)). On homophilic graphs, the global model generally shows better performance. In contrast, for heterophilic graphs, the aggregated global model underperforms relative to local training models. This discrepancy arises because clients with heterophilic graphs are more likely to exhibit diverse neighbor distribution patterns, leading locally trained GNNs to capture varied tendencies. As a result, the global model struggles to aggregate and generalize this knowledge, resulting in poor predictions. This preliminary experiment underscores the impact of distinct neighbor distribution patterns on FGL performance, reinforcing the need to address the challenges identified in our study.
@@ 3.1 Graph Structure Learning in FedHERO
Graph structure learning endeavors to comprehend the message-flow pattern throughout the graph (Xia et al., 2021). Nevertheless, in FL scenarios where data is diversely distributed and privacy concerns are paramount, directly aggregating all data to train a comprehensive structure learning model becomes unfeasible. To tackle this challenge, we implement a structure learning model for each client, which can learn structural knowledge from local data. By sharing the learned knowledge with other clients, the entire FGL system can acquire a unified, globally optimized structural representation that captures universally applicable patterns across diverse graphs.
@@
The objective of the structure learning model is to generate relevant latent structures that facilitate effective message passing based on the original graph. A common approach considers representing the weight of an edge between two nodes as a distance measure between two end nodes (Zhao et al., 2021b; Kazi et al., 2022; Chen et al., 2020). We follow this idea and employ a one-layer GNN to derive node representations $\mathbf{Z}$ from the original graph. The structure learning model can be represented by learning a metric function $\phi(\cdot,\cdot)$ of a pair of representations: $\tilde{a}_{uv}=\phi(\mathbf{z}_{u},\mathbf{z}_{v})$, where $\mathbf{z}_{u}$ is the learned embedding of node $u$ produced by GNN. The metric function $\phi(\cdot,\cdot)$ can assume various forms like cosine distance (Nguyen & Bai, 2010) and RBF kernel (Li et al., 2018b). In this study, we employ the multi-head attention mechanism as the metric function, following the approach outlined by Chen et al. (2020):
@@
This design choice is based on the differentiability of the top-k function, which facilitates the computation of parameter gradients and subsequent model updates in FedHERO. Additionally, we explore an alternative method for latent graph generation. Such a method treats each edge in the latent graph as a Bernoulli random variable. Here, each latent edge $\tilde{\mathbf{A}}_{ij}$ is sampled from a Bernoulli distribution with parameter $\tilde{a}_{ij}$ (Wu et al., 2020a; Elinas et al., 2020; Zhao et al., 2023). A comparison of these methods and the impact of varying k values is provided in Appendix C.
@@ 3.2 Optimization of the Dual-channel Model
In existing FGL frameworks (Zhang et al., 2021; Baek et al., 2023), the local models primarily employ one or multiple GNNs in sequence. These models initiate the process of generating node representations directly from raw features. However, as Tan et al. (2023) pointed out, directly constructing and sharing these local models across diverse clients can significantly compromise personalized performance due to feature and representation space misalignment. To address this challenge, divide the local model into two channels: $f_{global}$ and $f_{local}$. The global learner $f_{global}$ takes the original input graph and refines the latent graph structure. Through cooperation with other clients, $f_{global}$ detects patterns of universal relevance across heterophilic graphs. On the other hand, the local task learner $f_{local}$ extracts valuable client-specific information from raw node features and structures within the local graph. This information includes biased structures or diverse feature distributions tailored for follow-up tasks.
In a client’s local graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, we initiate the process by utilizing a Multi-Layer Perception (MLP) $f_{0}$ to extract information from the node feature matrix $\mathbf{X}$ into node embeddings $\mathbf{Z}^{0}$. Next, we feed the node embeddings $\mathbf{Z}^{0}$, the latent graph $\tilde{\mathbf{A}}$, and the original graph structure $\mathbf{A}$ into an L-layer GNN. The middle part of Figure 2 depicts how FedHERO leverages information from both the local graph structure $\mathbf{A}$ and the structure $\tilde{\mathbf{A}}$ from the structure learner to obtain node representations. In the $f_{global}$ channel (models in yellow boxes in Figure 2), the l-th layer of the GNN (denoted as $f_{g}^{l}$) utilizes the node embeddings from the previous layer $\mathbf{Z}^{l-1}$ and the latent graph $\tilde{\mathbf{A}}$ to generate hidden node embeddings that reveal patterns that hold universally across different graphs, denoted as $\mathbf{H}^{l}$. While feature propagation on the $\tilde{\mathbf{A}}$ could potentially yield enhanced node embeddings that align with global patterns, it is important to note that the original graph structures also contain valuable information, such as effective inductive bias (Bronstein et al., 2017). Therefore, we leverage the information within $\mathbf{A}$. It is combined with the output from the last layer, $\mathbf{Z}^{l-1}$, and fed into the l-th layer of the GNN (denoted as $f_{loc}^{l}$) within the $f_{local}$ channel (models in blue boxes in Figure 2). This process generates hidden node embeddings that adhere to the message-passing patterns within the local graph, resulting in $\mathbf{E}^{l}$. These embeddings play a crucial role in computing the layer-wise updates for node representations:
@@
Recent studies show that merging intermediate node representations from multi-hops enables the GNN to integrate nodes of the same class (Zhu et al., 2020; Abu-El-Haija et al., 2019). In addition, encoding the features of the ego node separately from the aggregated neighbor representations enhances node embedding learning (Zhu et al., 2020; Suresh et al., 2021; Yan et al., 2022). Thus, we merge the intermediate node representations generated by each layer of the dual-channel GNNs with the ego feature $\mathbf{X}$: $\mathbf{Z}^{out}=\textrm{COMB}(\mathbf{X},\mathbf{Z}^{0},...,\mathbf{Z}^{L}).$ The COMB function can be implemented in several ways, including max-pooling, concatenation, or LSTM-attention (Xu et al., 2018b). In our approach, we specifically use concatenation for the COMB function. This choice is based on the observation that concatenation consistently outperforms other methods in similar contexts (Xu et al., 2018b; Zhu et al., 2020).
@@
where $||\cdot||_{F}$ is the Frobenius norm. The first term in Eq. (7) evaluates the smoothness of the latent graph, operating under the assumption that nodes with smoother features are more likely to be connected. The second term in Eq. (7) serves the purpose of preventing excessively large node degrees. The hyperparameters $\lambda$ and $\mu$ govern the magnitude of the regularization effects. Hence, in conjunction with the cross-entropy loss $\mathcal{L}_{ce}$, we fine-tune the local model using the following loss function:
@@ 3.3 Model Aggregation in FedHERO
The right side of Figure 2 illustrates the process of clients sharing local models and receiving updated models from the server. The model aggregation process in FedHERO resembles that of FedAvg (McMahan et al., 2017). However, in the FedHERO framework, the server only receives and aggregates models from the $f_{global}$ channel of clients instead of the whole model. This sharing mechanism offers two key advantages for FedHERO. First, the structure learning model in the global channel generates latent graphs from node embeddings with relatively low dependence on the original graph structure. By sharing this model among clients, the global channel GNNs learn similar patterns from the generated latent graphs. Aggregating these aligned models enhances the effectiveness of FGL, leading to improved overall performance. Second, retaining the model $f_{local}$—which focuses on learning the message passing pattern within the local graph—as private and training it locally allows the local model to capture client-specific information, ultimately contributing to subsequent tasks. Let $w_{g}^{i}$ denote the parameters of the $f_{global}$ in client $i$. The server does the aggregation: $\overline{w}_{g}=\sum_{i}\frac{N_{i}}{N}w_{g}^{i}$ where $N_{i}$ represents the number of nodes in the client $i$’s graph, and $N$ denotes the total number of nodes in the graph data of all clients. Subsequently, the server transmits $\overline{w}_{g}$ to the clients. The clients then update model in the $f_{global}$ channel with $\overline{w}_{g}$ and commence local training for the ensuing round.